OpenStudy (anonymous):

y=x^2+1 y=x+1 Help please

OpenStudy (ranga):

Find the point where the two graphs intersect? If so set x^2 + 1 = x + 1 and solve for x. Then put that x value in the second equation and find y.

OpenStudy (anonymous):

yes it to find the point but i don't understand how to do the math. I can put in a graphing calculator but i need to under stand how to do it without the calculator

OpenStudy (ranga):

x^2 + 1 = x + 1 subtract 1 from both sides: x^2 = x subtract x from both sides: x^2 - x = 0 factor out x x(x - 1) = 0 either x = 0 or x = 1 Put x = 0, in second equation: y = x + 1 = 0 + 1 = 1. So (0,1) is one point of intersection. Put x = 1, in second equation: y = x + 1 = 1 + 1 = 2. So (1,2) is the second point of intersection.

OpenStudy (anonymous):

thanks i'll try this model on my other questions.

OpenStudy (ranga):

you are welcome.